OpenStudy (anonymous):

he perimeter of a square is 100 inches. What is the length of the diagonal of the square? Leave your answer in radical form.

OpenStudy (anonymous):

100/4=25 the square root of(25^2+25^2)=35.35

OpenStudy (anonymous):

that is your diagonal ...i used pythag theorem

OpenStudy (anonymous):

so how would that look in radical form

OpenStudy (anonymous):

sorry im not familiar with radical form

OpenStudy (anonymous):

could you explain what it is...and i might be able to asist you

OpenStudy (anonymous):

i have it but radical form is basically saying (25^2+25^2)= 25 square root of 2

OpenStudy (anonymous):

yes :)...

OpenStudy (anonymous):



OpenStudy (anonymous):

that would be it